FOURIER TRANSFORM MICROWAVE SPECTRA OF CO_{2}-ETHYLENE SULFIDE,CO_{2}-ETHYLENE OX-IDE AND CO_{2}-PROPYLENE OXIDE COMPLEXES

\underline{YUKARI ORITA} ,YOSHIYUKI KAWASHIMA , Department of Applied Chemistry, Faculty of Engineering, Kanagawa Institute of Technology, Atsugi, Kanagawa 243-0292, JAPAN; and EIZI HIROTA, The Graduate University for Advanced Studies, Hayama, Kanagawa 240-0193, JAPAN.

We have previously examined the difference in roles of O and S in structure and dynamics of the CO-ethylene oxide (EO) and CO-ethylene sulfide (ES) complexes^{a}. We have extended the investigation to CO_{2}-EO and CO_{2}-ES for comparison. We have also observed the CO_{2}-propylene oxide (PO) complex, which is an important intermediate in the reaction of PO with CO_{2} leading to polycarbonate. Both a-type and b-type transitions were observed for the CO_{2}-EO and CO_{2}-ES, but no c-type transitions were observed at all. We also detected the ^{34}S and ^{13}C isotopic species in natural abundance and the species containing ^{18}OCO and C^{18}O_{2}, which were synthesized by burning paper in an ^{18}O_{2} and^{16}O_{2} mixture. By analyzing the observed spectra we concluded the CO_{2} moiety of CO_{2}-EO and CO_{2}-ES located in a plane prependicular to the three-membered ring and bisecting the COC or CSC angle of EO or ES, respectively, as in the case of CO-EO and CO-ES complexes. An ab initio MO calculation at the level of MP2/6-311G(d, p) yielded an optimized structure in good agreement with the experimental result. We have derived from the observed spectra the distance, the stretching force constant, and the binding energy of the bonds between the constituents of the CO_{2}-EO and CO_{2}-ES complexes and have found that the distances of the two complexes were shorter by 0.2Å than those in CO-EO and CO-ES, respectively, and that the intermolecular bonds were two times stronger in the CO_{2} complexes than in the corresponding CO complexes. We have concluded from the observed spectra that the CO_{2} moiety in CO_{2}-PO is located on the PO three-membered ring plane opposite to the methyl group. The constituents in CO_{2}-PO were more weakly bound than those in CO_{2}-EO and CO_{2}-ES.

aS. Sato,Y. Kawashima, Y. Tatamitani, and E. Hirota,63rd International Symposium on Molecular Spectroscopy, WF05 (2008).